# DeepTrace: Learning to Optimize Contact Tracing in Epidemic Networks with Graph Neural Networks

### v3

@@ IV-B Algorithm DeepTrace: Pre-training and Fine-tuning GNNs
Output: The likelihood probability for every node in the network to be the contagion source.
@@
1. Pre-train the GNN model using contact tracing networks with node labels as the approximations of ~$\widetilde{\mathbb{P}}(G_{n}|v)$ in (11).
2. Fine-tune the GNN model using contact tracing networks with node labels having the exact values of ${\mathbb{P}}(G_{n}|v)$ in (1).
@@
IV-C Training Process with Semi-supervised Learning
To train the GNN model in DeepTrace during the two-phase training process, we define the loss functions for the pre-training and fine-tuning phase, respectively, as follows:
$L_{p}(\widehat{\mathbf{w}},\widecheck{\mathbf{w}},v\mid v\in G_{n})$
and
$L_{f}(\widehat{\mathbf{w}},\widecheck{\mathbf{w}},v\mid v\in G_{n})$
where ~$\widetilde{\mathbb{P}}(G_{n}\mid v)$ is the approximation of $\mathbb{P}(G_{n}\mid v)$, $\widehat{\mathbf{w}}=(\widehat{\mathbf{w}}^{(1)},\cdots,\widehat{\mathbf{w}}^{(L)})$ and $\widecheck{\mathbf{w}}=(\widecheck{\mathbf{w}}^{(1)},\cdots,\widecheck{\mathbf{w}}^{(L)})$. We take the logarithm of the likelihood probabilities to avoid arithmetic underflow, as the values of likelihood probabilities can be extremely small as the infected network grows in size.
Next, we describe how to construct the training data sets for pre-training and fine-tuning the GNN model, respectively.
IV-C1 Data Annotation for Pre-training
For the labeled training data in the pre-training phase, we can use an approximate estimator of (1) to efficiently obtain ~$\widetilde{\mathbb{P}}(G_{n}\mid v)$ for each node $v$ in the epidemic networks, minimizing time and effort. When the epidemic networks are degree-irregular trees, the ML estimator in (1) aims to calculate the sum of all probabilities of permitted permutations. We can rewrite (1) as follows:
$=\arg\max_{v\in G_{n}}\widetilde{\mathbb{P}}(G_{n}\mid v).$
IV-C2 Data Annotation for Fine-tuning
Since we already have a preliminary GNN model from pre-training, we can further refine it using a small amount of high-quality data. Specifically, this involves using data from epidemic networks with exact likelihood probabilities assigned to each node. To construct a piece of labeled data with exact likelihood probability for fine-tuning, a natural idea is to calculate the likelihood probability of each node directly by using (3) or (4). We first generate an epidemic network $\mathbb{G}_{N}$ with $N$ nodes, then for each node $v\in\mathbb{G}_{N}$ we generate all permitted permutation $\sigma\in\Omega(\mathbb{G}_{N}\mid v)$ starting from $v$ and compute $\sum_{\sigma\in\Omega(G_{n}\mid v)}\mathbb{P}(\sigma|v)$ using (3). Furthermore, we can construct data samples for which exact likelihood probabilities are easily obtainable using (4), such as the $d$-regular tree. In this case, since the probabilities of all permitted permutations are identical, it is evident from (10) that the likelihood probability for a node in a $d$-regular tree, as given by (2), is proportional to the epidemic centrality in [17], which can be computed using the message-passing algorithm described in [17].
IV-D Construction of Node Features
To ensure the effectiveness of GNN learning for solving (2) in Algorithm DeepTrace, it is crucial to meticulously design the node features. This careful design enables the extraction of node information from the underlying epidemic network $\mathbb{G}_{N}$, which corresponds to a specific spreading model during the training phase. We now consider several node features such as the proportion of instantaneously infected nodes, and boundary distances within the graph. Notably, constructing these features requires only linear computational complexity, $O(N)$, as illustrated in Fig. 5.
$\mathbf{Infected\;proportion}$: This is the ratio of the number of infected neighbors of a node $v_{i}$ to the number of all its neighbors:
$\hat{r}(v_{i})=\frac{\hat{d}(v_{i})}{d(v_{i})},$
where $\hat{d}(v_{i})$ is the number of infected neighbors of the node $v_{i}$. For example, the number of infected neighbors of nodes $v_{1},v_{2},v_{3},v_{4},$ and $v_{5}$ in Fig. 5 are $1,1,4,1$ and $1$, leading to an infected proportion of nodes $v_{1},v_{2},v_{3},v_{4},$ and $v_{5}$ being $\frac{1}{2},\frac{1}{3},1,\frac{1}{3},$ and $\frac{1}{4}$, respectively. This feature is to represent each node’s ability to infect new individuals. Intuitively, a node with a lower infected proportion has a higher probability of infecting new individuals and a smaller probability of being the source. Moreover, this particular node feature aims to capture the effect of the particular node position in the permitted permutation in (3). A node with a lower infected proportion ratio tends to be located at the front of the permitted permutation order, resulting in a smaller permutation probability. For example, as in Fig. 5, the different positions of Node $5$ for the permitted permutation $\{3,1,2,4,5\}$ and $\{5,3,1,2,4\}$ result in a probability of $1.7857\times 10^{-3}$ and $1.0204\times 10^{-3}$, respectively.
$\mathbf{Boundary\;distance\;ratio}$: As the infectivity of the virus evolves over time–newly symptomatic infectors generally exhibit higher infectivity, whereas long-term symptomatic infectors display lower infectivity–we assume that nodes at the boundary of the epidemic network are newly symptomatic infectors, while those further from the boundary are long-term symptomatic infectors. This characteristic is captured by the boundary distance ratio feature, which is the ratio of the shortest distance from the node $v_{i}$ to the farthest leaf node at the network boundary, as shown in Fig. 5, and is denoted by
$\check{r}(v_{i})=\frac{b(v_{i})}{\max_{v_{j}\in G_{n}}b(v_{j})},$
where $b(v_{j})$ is the boundary distance of node $v_{j}$. For example, the boundary distances of node $v_{1},v_{2},v_{4}$ and $v_{5}$ in Fig. 5 are $2$, and the boundary distances of node $v_{3}$ is $3$, so that the boundary distance ratio of nodes $v_{1},v_{2},v_{3},v_{4}$ and $v_{5}$ are $\frac{2}{3},\frac{2}{3},1,\frac{2}{3}$ and $\frac{2}{3}$, respectively. Notice that we grow $G_{i}$ using tree traversal algorithms; hence $G_{i}$ can be treated as a tree, and the boundary distance ratio of each node in $G_{i}$ is well-defined. From (2), we see that finding the source also requires information on the size of the collection of all permitted permutations $\Omega(G_{n}\mid v)$. Notice that a node with a larger boundary distance ratio tends to have a larger collection of permitted permutations. For example, there are $24$ permitted permutations when we choose node $v_{3}$ to be the source, but only $6$ permitted permutations if we choose either node $v_{1},v_{2},v_{4}$ or $v_{5}$ to be the source.
Figure 5: An example of a given epidemic network illustrating the boundary distance for each node.
@@ V Performance Evaluation
In this section, we first provide simulation results on the performance of the approximate ML estimator (11). Secondly, we compare the BFS and DFS strategies by simulating the procedure of forward and backward contact tracing. Lastly, we demonstrate the effectiveness of Algorithm DeepTrace on both synthetic networks and real-world networks based on COVID-$19$ contact tracing data in Hong Kong and Taiwan. As mentioned in Section I-A, the methods proposed in [30, 29] are based on the assumptions of either SEIR or SIR spreading model which is different from the SI model considered in this paper. We thus focus on baseline comparison relevant to the SI model, namely the rumor center heuristic in [15] and the SCT (Statistical Distance-Based Contact Tracing) algorithm in [23]. The software implementation of Algorithm DeepTrace and data used in our experiments can be found in https://github.com/convexsoft.
@@ V-A DeepTrace for Synthetic Networks
Data size Nodes Edges Labels Task
@@
TABLE I: The summary of the datasets of the synthetic epidemic networks for pre-training, fine-tuning and testing when training the GNN in DeepTrace.
[…]
@@
Among $1000$ generated synthetic epidemic networks, we use $50\%$ for pre-training, $25\%$ for fine-tuning, and $25\%$ for testing. Correspondingly, these datasets have approximately $50\sim 1000$, $50$, and $50\sim 1000$ infected nodes in the synthetic epidemic networks. We then construct the initial node features $\mathbf{h}_{v}^{(0)}=[\mathbf{1},\hat{r}(v),\check{r}(v)]^{T}$ for each node in these networks according to the analysis in the previous section. For the training data used in the pre-training phase, we label each node with ~$\widetilde{\mathbb{P}}(G_{n}\mid v)$ in (11), while for the network data used in the fine-tuning phase and testing, we label each node with the exact probability ${\mathbb{P}}(G_{n}\mid v)$ in (2). A summary of these datasets is shown in TABLE I. Then we pre-train and fine-tune the GNN model in DeepTrace both for $150$ epochs, and use the bias and a top-$k$ accuracy to evaluate the performance of the trained model in the two-phase training process. The “top-$k$" metric signifies that the trained GNN identifies the approximate superspreader as one of the $k$ nodes with the highest probability of being superspreaders. This metric is similar to the pass@k metric for evaluating the functional correctness of code generated by pre-trained models in [45]. In particular, the pass@k metric is the probability that at least one of the top-$k$ generated code samples for a problem passes the unit tests [45]. In the context of evaluating the efficacy of pre-trained models in learning to optimize, this metric measures how well the ranked nodes in $G_{n}$ solves (1) based on the notion of network centrality. For example, if the approximate superspreader detected by the trained GNN is in the top-10 set, it indicates that this node is among the 10 most likely superspreaders. If it is in the top-1 set, it is the most likely superspreader. Thus, the lower the value of $k$ in the top-$k$ set, the higher the likelihood that the detected node is the most probable superspreader. Moreover, a higher likelihood of the optimal solution to (1) being among the top-$k$ for a smaller $k$ implies better performance of the pre-trained GNN. Thus, the “top-$k$" metric effectively measures the frequency with which the global optimal solution of (1) appears within the top-ranked prediction outputs of the pre-trained GNN, offering a quantitative assessment of the accuracy in learning to optimize. The results are shown in Fig. 6.
@@
Another key evaluation is to measure the generalizability of DeepTrace Algorithm, that is, to validate the effectiveness of the trained model by the small-scale epidemic networks in the larger epidemic networks. We construct six additional synthetic network validation datasets, each with $100$ networks, and the number of nodes in each set of networks is $50,100,200,500,1000,$ and $2500$, respectively. We use the SI model to simulate the spreading for a given number of infected nodes to generate the infection graphs (see Section IV-B). We then evaluate the performance of the trained GNN on these validation datasets using a metric based on top-$k$ accuracy.
After training for 150 epochs, we evaluate the prediction accuracy at top-1, top-10, and top-20 for each validation dataset. The results of the pre-training and fine-tuning processes are illustrated in Fig. 7 and Fig. 8, respectively. Fig. 7 displays the accuracy trends during the pre-training phase, showing a steady improvement in performance as the model learns. In contrast, Fig. 8 depicts the fine-tuning phase results, highlighting how the model adjusts and refines its predictions with additional training. These figures collectively demonstrate that the training accuracy remains relatively stable, not deteriorating significantly even as the number of nodes in the epidemic networks increases. This indicates robust performance and adaptability of the GNN across varying network sizes. This phenomenon underscores the utility of developing data-driven models using small-sized networks for training, with the expectation that these models can generalize effectively to larger-sized networks. Utilizing smaller networks not only reduces overall training time but also facilitates leveraging transfer learning to extend models for superspreader detection in larger networks.
@@ V-B Evaluation on BFS and DFS tracing
In this section, we compare the performance of BFS and DFS tracing strategies using first detection time and average error as evaluation metrics, defined by equations (7) and (8), respectively. Here we take into account five groups of epidemic network datasets with distinct topological structure. The relevant statistics for these data are displayed in TABLE II. In the final column, we use $d$ to represent the degree of the regular network, $p$ to indicate the likelihood of edge formation between nodes in the ER random network, and $NC$/$p$/$q$ to denote the number of communities, the probability of edge formation within a community, and the probability of edge formation between communities, respectively. Additionally, we define the sensor network’s coverage area as $Area$ and the communication range, denoted as $CR$, as the utmost distance between sensors that can form an edge. Let us describe the process of the simulation experiment. In synthetic networks, we first establish the underlying network $\mathbb{G}$ using different random network models. Next, based on the SI model described in Section II-A, we uniformly and randomly select a source node and begin simulating the spread of the virus on $\mathbb{G}$. Since the number of nodes in our real-world network data is around 100, we stop this spreading process once the number of infected nodes reaches 100, resulting in a 100-node epidemic network $\mathbb{G}_{N}$. To ensure fairness (as different index nodes may lead to significant performance differences), each node in $\mathbb{G}_{N}$ is selected as the index node to start the contact tracing process.
@@
We perform forward contact tracing using BFS or DFS and employ Algorithm DeepTrace along with two other methods, Rumor Center [15], and SCT [23], for backward contact tracing. As we expand the contact tracing network $G_{n}$, each time a new node is added, we compute the current estimated superspreader in $G_{n}$ and measure the distance between this estimated superspreader and the true source, noting the error in hops at that moment (cf. (7)). When the estimated superspreader first corresponds to the true source, we record the current size of $G_{n}$ as the first detection time. When $G_{n}=\mathbb{G}_{N}$, we will have identified $100$ approximated superspreaders and their corresponding error hops. We calculate the average of these 100 error hops as the average error hops for this simulation experiment. This experiment is repeated hundreds of times on the same underlying network $\mathbb{G}$, but with different epidemic networks $\mathbb{G}_{N}$, recording the average error hops, first detection time, and the time spent. We initiate the contact tracing simulation directly on this data for the real-world network data (Taiwan and Hong Kong COVID-19 datasets), as these networks already represent epidemic conditions.
@@
The theoretical result we obtained in Theorem 1 for the BFS strategy describe the distance between the approximated superspreader and the MLE, whereas the theoretical results, from Lemma 1 and Theorem 2, for the DFS strategy calculate the upper bound of the number of nodes that will be marked as approximated superspreaders during the contact tracing process. The theoretical results for BFS and DFS are entirely different in scope, so we cannot determine which strategy is better based on theory alone. In our simulation experiments, we tested six combinations of DeepTrace, Rumor Center, and STC with either BFS or DFS and recorded the first detection time in Table III and the average error in Table V. In addition, we also verify the results of Lemma 1 and Theorem 2 in Table IV. In Table III, we can observe that DeepTrace+BFS outperforms other methods from the first detection perspective. Additionally, within Rumor Center and STC, the BFS strategy generally outperforms the DFS strategy. On the other hand, from the perspective of average error, when the underlying network is a tree, DeepTrace+DFS performs better. Otherwise, DeepTrace+BFS is the best choice.
In addition to testing the performance of different methods on average error and first detection time, we also evaluated the computational efficiency of DeepTrace, Rumor Center, and STC. For computational efficiency, we recorded the time taken for contact tracing using DeepTrace, Rumor Center, and STC combined with DFS and BFS and calculated the average for DFS and BFS. As Table VI shows, DeepTrace’s computational speed is significantly higher than that of the other two methods.
@@
Epidemic Network DeepTrace Rumor center SCT
DFS BFS DFS BFS DFS BFS
Complete N-ary graph 34.00 7.17 34.81 8.18 38.60 16.95
Regular tree 36.10 16.14 63.06 18.21 64.44 46.25
ER random graph 48.27 12.48 65.13 10.97 23.56 36.75
SBM network 34.07 47.14 49.03 54.98 44.56 40.7
Sensor network 32.47 18.71 33.45 21.14 39.28 36.75
Real-world network (Taiwan) 1.18 0.55 2.95 3.08 4.59 4.61
TABLE III: The average first detection time of DeepTrace and baseline methods across different types of epidemic networks, utilizing BFS and DFS tracing.
Epidemic Network Forward tracing DeepTrace Rumor center SCT
$\frac{|S_{1}|}{N-1}$ $\frac{|S_{1}|}{N-2}$ $\frac{|S_{3}|}{N-1}$ $\frac{|S_{1}|}{N-1}$ $\frac{|S_{2}|}{N-1}$ $\frac{|S_{3}|}{N-1}$ $\frac{|S_{1}|}{N-1}$ $\frac{|S_{2}|}{N-1}$ $\frac{|S_{3}|}{N-1}$
Complete N-ary graph DFS 0.78 0.08 0.13 0.97 0.02 0.01 0.96 0.03 0.01
BFS 0.88 0.08 0.04 0.97 0.02 0.01 0.98 0.02 0.00
Regular tree DFS 0.66 0.21 0.13 0.91 0.06 0.03 0.89 0.07 0.04
BFS 0.69 0.17 0.13 0.92 0.05 0.03 0.94 0.05 0.01
ER random graph DFS 0.71 0.16 0.13 0.94 0.04 0.02 0.96 0.03 0.01
BFS 0.79 0.11 0.10 0.95 0.03 0.02 0.96 0.03 0.01
SBM network DFS 0.74 0.12 0.14 0.94 0.04 0.02 0.93 0.05 0.02
BFS 0.83 0.08 0.09 0.95 0.03 0.02 0.96 0.03 0.01
Sensor network DFS 0.76 0.14 0.10 0.94 0.04 0.02 0.94 0.04 0.02
BFS 0.83 0.09 0.08 0.94 0.04 0.02 0.94 0.04 0.02
Real-world network (Taiwan) DFS 0.87 0.07 0.06 0.97 0.02 0.01 0.97 0.02 0.01
BFS 0.68 0.23 0.09 0.98 0.01 0.01 0.98 0.02 0.00
TABLE IV: Statistical analysis of consecutive pairs of approximated superspreader for DeepTrace and baseline methods across various types of epidemic networks employing BFS and DFS tracing. Note that the value of $\frac{|S_{i}|}{N-1}$ represents the percentage of $|S_{i}|$ in the total quantity, which is $|S_{1}|+|S_{2}|+|S_{3}|$.
Epidemic Network DeepTrace Rumor center SCT
DFS BFS DFS BFS DFS BFS
Complete N-ary graph 0.84 1.54 1.35 2.31 1.04 1.66
Regular tree 2.37 2.45 2.63 2.56 2.62 3.30
ER random graph 2.70 2.52 2.93 2.92 2.78 2.55
SBM network 2.55 2.30 2.78 2.86 2.73 2.69
Sensor network 2.67 2.61 3.50 3.42 3.06 2.85
Real-world network (Taiwan) 0.97 0.65 1.38 0.88 0.79 0.80
TABLE V: The average error during contact tracing for Algorithm DeepTrace and the baselines on various types of epidemic networks using BFS and DFS tracing strategies.
@@
TABLE VI: The efficiency comparison, in terms of time required (measured in seconds) for contact tracing, between Algorithm DeepTrace and other baseline methods across epidemic networks of varying scales.

### v4

@@ IV-E The DeepTrace Algorithm
Output: The log-likelihood for every node in the network to be the contagion source.
@@
1. Pre-train the GNN model using contact tracing networks with node labels as the approximations of $\log\widetilde{\mathbb{P}}(G_{n}|v)$ in (11).
2. Fine-tune the GNN model using contact tracing networks with node labels having the exact values of $\log{\mathbb{P}}(G_{n}|v)$ in (1).
@@
As contact tracing progresses, Algorithm 1 ultimately identifies the superspreaders for the entire epidemic network. The computational complexity of contact tracing in Algorithm 1 can be decomposed into three main components: sampling complexity, aggregation complexity, and the complexity of the final regression layer. Suppose there are $N$ nodes in the epidemic network. During the sampling procedure, the GNN samples a fixed-size neighborhood for each node. If each node samples $S$ neighbors and this process is repeated for $L$ layers, then the total number of nodes sampled for each node is $O(S^{L})$. In the final step of contact tracing, the contact tracing network contains $N$ nodes, so the sampling complexity for this network is $O(NS^{L})$. Since we initiate contact tracing from the index case, the number of nodes in the contact tracing network increases sequentially. Consequently, the overall sampling complexity for the entire contact tracing process across the epidemic network is $O(N^{2}S^{L})$. For the aggregation procedure, the GNN aggregates information from its neighbors sampled. Assume that the feature dimension is $m$. Since the LSTM aggregator processes the features of the sampled neighbors, the complexity of the aggregation step is $O(mS^{2})$ per node per layer. Thus, the aggregation complexity for contact tracing the entire epidemic network is $O(N^{2}LmS^{2})$. In the final regression layer, a fully connected layer maps node embeddings to target values, representing the likelihood of nodes being superspreaders. The per-node complexity is $O(m)$, resulting in $O(N^{2}m)$ for the whole network. Thus, the overall computational complexity of contact tracing in Algorithm 1 is $O(N^{2}LmS^{2})$.
@@ V Performance Evaluation
In this section, we first provide simulation results on the performance of the approximate ML estimator (11). Secondly, we compare the BFS and DFS strategies by simulating the procedure of forward and backward contact tracing. Lastly, we demonstrate the effectiveness of DeepTrace on both synthetic networks and real-world networks based on COVID-$19$ contact tracing data in Hong Kong and Taiwan. As mentioned in Section I-A, the methods proposed in [30, 29] are based on the assumptions of either SEIR or SIR spreading model, which is different from the SI model considered in this paper. We thus focus on baseline comparison relevant to the SI model, namely the rumor center heuristic in [15], and the SCT (Statistical Distance-Based Contact Tracing) algorithm in [23]. The software for Algorithm DeepTrace and experimental data are available at https://github.com/convexsoft.
@@ V-A DeepTrace for Synthetic Networks
We first evaluate the performance of Algorithm DeepTrace for identifying superspreaders within epidemic networks, using the testing datasets, after we pre-train the GNN model for $150$ epochs. Then, we evaluate the performance of DeepTrace again after another $150$ epoch training for fine-tuning. There are $250$ pieces of data in the testing dataset. The generation of each piece of data follows the procedure described in Section IV-C1 except the node label in the testing dataset is the true log-likelihood $\log{\mathbb{P}}(G_{n}|v)$. We summarize the dataset information, including datasets for pre-training, fine-tuning, and testing, in Table I.
TABLE I: Datasets of the synthetic epidemic networks for pre-training, fine-tuning, and testing when training the GNN in DeepTrace. The first column is the number of networks used in each phase.
No. $|V(\mathbb{G}_{N})|$ $|E(\mathbb{G}_{N})|$ Labels Task
@@
We use the bias [31] and a top-$k$ accuracy to evaluate the performance of the trained model in the two-phase training process. The “top-$k$" metric signifies that the trained GNN identifies the approximate superspreader as one of the $k$ nodes with the highest probability of being superspreaders. This metric is similar to the pass@k metric for evaluating the functional correctness of code generated by pre-trained models in [47]. In particular, the pass@k metric is the probability that at least one of the top-$k$ generated code samples for a problem passes the unit tests [47]. In the context of evaluating the efficacy of pre-trained models in learning to optimize, this metric measures how well the ranked nodes in $G_{n}$ solve (1) based on the notion of network centrality. For example, if the approximate superspreader detected by the trained GNN is in the top-10 set, it indicates that this node is among the $10$ most likely superspreaders. Thus, the lower the value of $k$ in the top-$k$ set, the higher the likelihood that the detected node is the most probable superspreader. Moreover, a higher likelihood of the optimal solution to (1) being among the top-$k$ for a smaller $k$ implies the better performance of the pre-trained GNN. Thus, the “top-$k$" metric effectively measures the frequency with which the global optimal solution of (1) appears within the top-ranked prediction of the pre-trained GNN, offering a quantitative assessment of the accuracy.
Fig. 6 illustrates the value of bias and accuracy over the training epochs during both the pre-training and fine-tuning phases. Fig. 6 and 6 show that the bias decreases significantly as the number of epochs increases, with a faster and more pronounced reduction during fine-tuning. Initially, the pre-training phase exhibits higher bias, which gradually decreases and levels off after about 20 epochs, whereas fine-tuning achieves a lower bias within the first few epochs. Fig. 6 and 6 present the top-$1$, top-$5$, and top-$10$ accuracy trends, demonstrating that accuracy improves steadily throughout both phases. During pre-training, accuracy increases more slowly, with top-10 accuracy achieving the highest values, followed by top-5 and top-1 accuracy. In contrast, fine-tuning rapidly enhances accuracy, reaching a stable level within 20 epochs. These results indicate that fine-tuning not only accelerates convergence but also achieves a more optimal balance between bias reduction and accuracy improvement compared to the model with only the pre-training phase.
@@
Another key evaluation is to measure the generalizability of DeepTrace Algorithm, that is, to validate the effectiveness of the trained model by the small-scale epidemic networks in the larger epidemic networks. We construct six additional synthetic network validation datasets, each with $100$ random $3$-regular networks, and the number of nodes in each set of networks is $50,100,200,500,1000,$ and $2500$, respectively. We use the SI model to simulate the spreading for a given number of infected nodes to generate the infection graphs (see Section IV-C). We evaluate the trained GNN’s performance on validation datasets using top-$k$ accuracy.
After training for 150 epochs, we evaluate the prediction accuracy at top-1, top-10, and top-20 for each validation dataset. The results of the pre-training and fine-tuning processes are illustrated in Fig. 7 and Fig. 8, respectively. Fig. 7 displays the accuracy trends during the pre-training phase, showing a steady improvement in performance as the model learns. In contrast, Fig. 8 depicts the fine-tuning phase results, highlighting how the model adjusts and refines its predictions with additional training. These figures demonstrate that the training accuracy remains relatively stable, not deteriorating significantly even as the number of nodes in the epidemic networks increases. This demonstrates the GNN’s robustness and adaptability across network sizes, highlighting the utility of training on small networks for generalization to larger ones. Training on smaller networks reduces time and enables transfer learning for source detection in larger networks.
@@ V-B Evaluation on BFS and DFS tracing
In this section, we compare the performance of BFS and DFS tracing strategies using first detection time and average error as evaluation metrics, defined by equations (7) and (8), respectively. Here, we consider five groups of epidemic network datasets with distinct topological structures. The relevant statistics for these data are displayed in TABLE II. In the final column, we use $d$ to represent the degree of the regular network, $p$ to indicate the likelihood of edge formation between nodes in the ER random network, and $NC$/$p$/$q$ to denote the number of communities, the probability of edge formation within a community, and the probability of edge formation between communities, respectively. Additionally, we define the sensor network’s coverage area as $Area$ and the communication range, denoted as $CR$, as the utmost distance between sensors that can form an edge. Let us describe the simulation experiment process. In synthetic networks, we first establish the underlying network $\mathbb{G}$ using different random network models. Next, based on the SI model described in Section II-A, we uniformly and randomly select a source node and begin simulating the spread of the virus on $\mathbb{G}$. Since the number of nodes in our real-world network data is around 100, we stop this spreading process once the number of infected nodes reaches 100, resulting in a 100-node epidemic network $\mathbb{G}_{N}$. To ensure fairness (as different index nodes may lead to significant performance differences), each node in $\mathbb{G}_{N}$ is selected as the index node to start the contact tracing process.
@@
We perform forward contact tracing using BFS or DFS and employ Algorithm DeepTrace along with two other methods, Rumor Center [15], and SCT [23], for backward contact tracing. As we expand the contact tracing network $G_{n}$, each time a new node is added, we compute the current estimated superspreader in $G_{n}$ and measure the distance between this estimated superspreader and the true source, noting the error in hops at that moment (cf. (7)). When the estimated superspreader first corresponds to the true source, we record the current size of $G_{n}$ as the first detection time. When $G_{n}=\mathbb{G}_{N}$, we will have identified $100$ approximated superspreaders and their corresponding error hops. We calculate the average of these 100 error hops as the average error hops for this simulation experiment. This experiment is repeated $100$ times on the same underlying network $\mathbb{G}$, but with different epidemic networks $\mathbb{G}_{N}$, recording the average error hops, first detection time, and the time spent. We initiate the contact tracing simulation directly on this data for the real-world network data (Taiwan and Hong Kong COVID-19 datasets), as these networks already represent epidemic conditions.
@@
The theoretical result we obtained in Theorem 1 for the BFS strategy describe the distance between the approximated superspreader and the MLE, whereas the theoretical results from Lemma 1 and Theorem 2, for the DFS strategy calculate the upper bound of the number of nodes that will be marked as approximated superspreaders during the contact tracing process. The theoretical results for BFS and DFS are entirely different in scope, so we cannot determine which strategy is better based on theory alone. In our simulation experiments, we tested six combinations of DeepTrace, Rumor Center, and STC with either BFS or DFS and recorded the first detection time and the average distance error in Table III. We can observe that DeepTrace+BFS outperforms other methods from the perspective of the first detection time. Additionally, within Rumor Center and STC, the BFS strategy generally outperforms the DFS strategy. On the other hand, from the perspective of the average distance error, when the underlying network is a tree with a symmetrical structure, such as complete N-ary graphs or regular trees, DeepTrace+DFS performs better. In summary, DeepTrace generally exhibits lower average distance errors than the baseline methods. BFS tracing often performs better than DFS for DeepTrace, indicating that BFS may be more effective for these structures. The results show that DeepTrace achieves higher accuracy in complex network topologies and consistently outperforms baseline methods across various networks. We evaluated the computational efficiency of DeepTrace, Rumor Center, and STC, measuring the average contact tracing time using DFS and BFS. As shown in Table IV, DeepTrace achieves significantly higher computational speed than the other methods.
@@
TABLE III: Comparison of Average First Detection Time and Average Distance Error for DeepTrace and baseline methods across different epidemic networks, using BFS and DFS tracing strategies.
Epidemic Network Average First Detection Time Average Distance Error
DeepTrace Rumor Center SCT DeepTrace Rumor Center SCT
DFS BFS DFS BFS DFS BFS DFS BFS DFS BFS DFS BFS
Complete N-ary graph 34.00 7.17 34.81 8.18 38.60 16.95 0.84 1.54 1.35 2.31 1.04 1.66
Regular tree 36.10 16.14 63.06 18.21 64.44 46.25 2.37 2.45 2.63 2.56 2.62 3.30
ER random graph 48.27 12.48 65.13 10.97 23.56 36.75 2.70 2.52 2.93 2.92 2.78 2.55
SBM network 34.07 47.14 49.03 54.98 44.56 40.70 2.55 2.30 2.78 2.86 2.73 2.69
Sensor network 32.47 18.71 33.45 21.14 39.28 36.75 2.67 2.61 3.50 3.42 3.06 2.85
Real-world network (Taiwan) 1.18 0.55 2.95 3.08 4.59 4.61 0.97 0.65 1.38 0.88 0.79 0.80
TABLE IV: The efficiency comparison, in terms of time required (measured in seconds) for contact tracing, between Algorithm DeepTrace and other baseline methods across epidemic networks of varying scales.